# Generating Fock-state superposition from coherent state by quantum measurement

### v1

@@ Abstract
High-level Fock states and their superpositions are essentially exotic testbeds for nonclassical physics and valuable resources for quantum technologies. We provide a simple protocol on quantum measurement to generate an arbitrary Fock state and selected superposed Fock states from a coherent state of a target resonator, without any carefully tailored external driving. This conditional protocol can be efficiently constructed by a sequence of joint free-evolution of the resonator and an ancillary qubit, that are coupled via a Jaynes-Cummings interaction, and projective measurements on the qubit. By properly choosing the duration of each evolution-measurement cycle and the initial state of the resonator, we can generate a desired Fock state $\left|n\right\rangle$ and a superposed Fock state $(\left|0\right\rangle+\left|n\right\rangle)/\sqrt{2}$, $n\sim 10$, with a fidelity over $99\%$ in less than $30$ measurements. Moreover, our protocol can be straightforwardly extended to the generation of a multi-excitation Bell state $(\left|00\right\rangle+\left|nn\right\rangle)/\sqrt{2}$ in a double-resonator system.
@@ I Introduction
A Fock state is of the most nonclassical state with a well-defined number of bosonic excitations such as photons, phonons, and magnons. It is found to play diverse and crucial roles in quantum information fields Yamamoto and Haus (1986), including quantum communications Caves and Drummond (1994), quantum metrology Holland and Burnett (1993); Nagata et al. (2007); Giovannetti et al. (2011); Slussarenko et al. (2017), and remote entanglement generation Narla et al. (2016). Fock states are prominent candidates in open quantum systems to explore the environment-induced decoherence by measuring their decay time Wang et al. (2008); Brune et al. (2008). The superpositions of Fock states with particular components also exhibit appealing features in fault-tolerant quantum information processing Joshi et al. (2021). For example, the superposed Fock states with binomial coefficients can be used in quantum error correction Michael et al. (2016). And a high-precision measurement over weak classical signals can be achieved by using superposed Fock states Caves et al. (1980) due to the squeezing effect induced by quantum interference Wódkiewicz et al. (1987). Generating desired Fock states and their superposition is therefore of significant interest to both the fundamental quantum physics and the applications in quantum technology.
@@
Early proposals to generate Fock state employ the resonator and the ancillary atom in the Jaynes-Cummings (JC) model, which are under a full control in both initial states and coupling period. The atom serves as an intermediator between the resonator and the external driving pulse to pump quanta into the resonator by sequential half Rabi oscillations. The oscillation period depends on the current state of the resonator, by which the generation proposals demand elaborate external driving in the cavity QED setup Vogel et al. (1993); Law and Eberly (1996). Based on such a stepwise method, Fock states and their superpositions were then experimentally generated in coplanar waveguide resonators for a microwave photon number $n\leq 6$ with a fidelity over $90\%$ Hofheinz et al. (2008, 2009) and in a bulk acoustic-wave resonator for up to $n=7$ phonons with a fidelity over $26\%$ Chu et al. (2018). In the transmon-cavity platform, Fock state with $n\leq 3$ can be generated by dispersively coupling the transmon-qubit to the photonic resonator and using a two-photon stimulated Raman adiabatic passage Premaratne et al. (2017). Two-level atomic systems are not necessary in the step-by-step idea. A proposal in Ref. Yanagimoto et al. (2019) was presented to generate Fock state by the simultaneous adiabatic modulation over the driving frequency and intensity on a Kerr nonlinear cavity.
@@
Beyond the methods based on multiple Rabi oscillations or equivalent transitions, Uria et al. considered in a cavity QED setup, a two-level atom resonantly coupled to the target cavity field initially prepared in a coherent state Uria et al. (2020). It had been numerically found that after a particular interaction period, the cavity field evolves into a displaced Fock state of a large photon number with a fidelity over $70\%$. Recently, the experimental demonstration of a dissipation engineering over a three-dimensional microwave cavity was reported in Ref. Li et al. (2024), which is based on a cascaded selective photon-addition operation assisted by an ancillary superconducting qubit and a readout resonator. The readout resonator was used to implement a quantum reservoir control for stabilizing a multiphoton Fock state of the cavity mode with a photon number up to $n=3$ with a fidelity over $77\%$.
@@
Quantum measurement is found to be a promising candidate for state control tasks, such as quantum state purification Nakazato et al. (2003); Combes et al. (2010); Wiseman and Ralph (2006), ground state cooling of various quantum systems Li et al. (2011); Puebla et al. (2020); Yan and Jing (2021), and nuclear spin polarization Jin et al. (2022). Also, it can be implemented in generating Fock states by elaborate external tuning or by nondemolition measurements. For example, a Fock state of a cavity field can be conditionally created by the postselection over the atomic level after a specified period of evolution in a cavity QED setup, in which the atom-field interaction is highly tunable to control the transition rate in a chosen atom-field subspace on resonance Fran ça Santos et al. (2001). An arbitrary superposed Fock state can be generated by the dispersive coupling between the cavity field and a qubit so that one could individually control the qubit transition rate associated with the corresponding qubit-field subspace by driving the qubit with external microwave signals containing multiple frequency components Wang et al. (2017). On performing quantum nondemolition measurement and feedback control over the field photon number to conditionally suppress the photon number spread in Fock space Guerlin et al. (2007), a Fock state with a small photon number $n\leq 3$ was proposed in theory Dotsenko et al. (2009) and implemented in experiment Sayrin et al. (2011). Recently, quantum nondemolition measurement has been used in the preparation of a superposed Fock state in a circuit QED setup, which can be optimized by reinforcement learning Perret and Bérubé-Lauzière (2024).
@@
With a JC Hamiltonian, we propose a protocol to conditionally generate a desired Fock states $\left|n\right\rangle$ and selected superposed Fock states, such as $(\left|0\right\rangle+\left|n\right\rangle)/\sqrt{2}$, of a target resonator that starts from the coherent state without extra cost in state preparation. The protocol is based on repeated performing projective measurements on the coupled qubit, which are separated with joint free-evolution of the composite system lasting a period of $\tau$. For the resonator, the evolution-and-measurement cycles induce a positive operator valued measure (POVM) described by $V_{ji}(\tau)=\left\langle j\right|e^{-iH\tau}\left|i\right\rangle$, where $\left|i\right\rangle$ and $\left|j\right\rangle$ represent the initial state of the ancillary qubit and the measurement outcome, respectively and $H$ denotes the full Hamiltonian of the composite system. Our protocol does not require any elaborate external parametric driving or tuning during the whole generation process. By appropriately choosing POVM and the period $\tau$ for each evolution-measurement cycle, the selected Fock components can be efficiently distinguished from the coherent state. By using qutrit instead of qubit and two nondegenerate resonators instead of a single one, our protocol can be extended to generate double-mode Fock states and multi-excitation Bell states $(\left|00\right\rangle+\left|nn\right\rangle)/\sqrt{2}$. In terms of the JC model and the projective measurement over the ancillary system, our work is inspired by previous proposals for generating nonclassical bosonic states, e.g., the Fock states and their classical mixtures in Refs. Harel and Kurizki (1996) and Delakouras et al. (2023), respectively, and the squeezed state in Ref. Weiher et al. (2019). Nevertheless, our protocol is applicable to generate superposed states and extendable to multiple modes.
@@
The rest of this work is organized as follows. In Sec. II, we introduce the theoretical framework for POVM based on the JC model and repeated projective measurements on the ancillary qubit. We find the reduction factor on the population distribution of the resonator, which is crucial to the ensued state generation, is exclusively determined by the qubit-resonator coupling strength, their detuning, and the period of each evolution-and-measurement cycle. In Sec. III.1 and Sec. III.2, we discuss the generation of various single-mode Fock states and their coherent superpositions, respectively. The protocol efficiency can be optimized by modulating the evolution periods. In Sec. IV, our protocol is extended to generate the double-mode Fock states or multi-excitation Bell state of two nondegenerate resonator modes with 2D population-reduction factors and strategies. We summarize the whole work in Sec. V.
@@ II Theoretical framework
Figure 1: Model diagram for our superposed Fock-state generation protocol. A two-level system (qubit) serves as the ancillary system. The target resonator is coupled to the qubit through exchange interaction with a strength $g$. $\delta$ is their detuning in frequency. The projective measurements $M_{i}=|i\rangle\langle i|$, $i=g,e$, are performed on the ancillary qubit.
Our protocol to generate the desired Fock states or their superposition of a single bosonic mode is based on a composite system consisting of a target resonator coupled to an ancillary qubit with a JC interaction, as shown in Fig. 1. The full Hamiltonian reads ($\hbar=1$)
@@
where $\omega_{e}$ and $\omega_{f}$ are the level splitting of the ancillary qubit and the transition frequency of the target resonator, respectively. $a^{\dagger}$ ($a$) is the creation (annihilation) operator of the resonator. $g$ denotes the qubit-resonator coupling strength. In the rotating frame with respect to $U_{0}=\exp[i\omega_{f}(|e\rangle\langle e|+a^{\dagger}a)t]$. The Hamiltonian becomes
@@
where $\delta\equiv\omega_{e}-\omega_{f}$ denotes the detuning between the qubit and the resonator mode.
Suppose that the ancillary qubit is prepared as one of its energy eigenstate $\left|i\right\rangle$, $i\in\{e,g\}$, and the resonator is initially at $\rho_{a}(0)$. After a period $\tau$ of joint free-evolution governed by $U(\tau)=\exp(-iH^{\prime}\tau)$, we perform a projective measurement on the qubit. If the measurement outcome is $\left|j\right\rangle$, $j\in\{e,g\}$, the density matrix of the resonator is then found to be
@@
where $V_{ji}(\tau)=\left\langle j\right|e^{-iH^{\prime}\tau}\left|i\right\rangle$ denotes an effective (nonunitary) evolution operator acting on the resonator. And $P_{j}(1)$ represents the success probability of the ancillary qubit on the state $\left|j\right\rangle$ after an evolution-measurement cycle. In our protocol, the cycle is repeated if $\left|j\right\rangle=\left|i\right\rangle$. Otherwise, the entire protocol is restarted from the very beginning.
The effective evolution operator $V_{ii}(\tau)$ then can be relabeled as $V_{i}(\tau)$ for simplicity, which is diagonal in the Fock-state basis $\{\left|k\right\rangle\}$, i.e.,
$V_{i}(\tau)=\sum_{k=0}^{\infty}\lambda_{k}^{(i)}(\tau)\left|k\right\rangle\left\langle k\right|.$ (4)
@@
where the $k$-photon Rabi frequency for $i=e,g$ are
@@
respectively. Since $|\lambda^{(i)}_{k}(\tau)|\leq 1$, it works evidently as a reduction factor for the population on the Fock-state $\left|k\right\rangle$.
If the outcomes of each measurement are identical to the initial state of the qubit $\left|i\right\rangle$ and $\tau$ is invariant, the density matrix of the resonator under $N$ rounds of evolution and measurement is found to be
$\rho_{a}(N\tau)$ $=\dfrac{V_{i}^{N}(\tau)\rho_{a}(0)V_{i}^{\dagger N}(\tau)}{\text{Tr}\left\{V_{i}^{N}(\tau)\rho_{a}(0)V_{i}^{\dagger N}(\tau)\right\}}$
@@
$\mathcal{D}[\rho_{a}(N\tau)]$ $=\sum_{k=0}^{\infty}\frac{p_{k}|\lambda_{k}^{(i)}(\tau)|^{2N}}{P_{i}(N)}\left|k\right\rangle\left\langle k\right|,$ (9)
$\mathcal{C}[\rho_{a}(N\tau)]$ $=\sum_{k\neq k^{\prime}}\frac{C_{kk^{\prime}}\left[\lambda_{k}^{(i)}(\tau)\lambda_{k^{\prime}}^{(i)*}(\tau)\right]^{N}}{P_{i}(N)}\left|k\right\rangle\left\langle k^{\prime}\right|,$ (10)
where $\mathcal{D}[\rho_{a}]$ and $\mathcal{C}[\rho_{a}]$ denote the diagonal and off-diagonal parts of the density matrix, respectively. And $p_{k}\equiv\left\langle k\right|\rho_{a}(0)\left|k\right\rangle$ and $C_{kk^{\prime}}\equiv\left\langle k\right|\rho_{a}(0)\left|k^{\prime}\right\rangle$ are the populations and coherence of the initial density matrix of the resonator in the Fock-state basis $\{\left|k\right\rangle\}$. The success probability of finding the qubit at its initial state after $N$ measurements is found to be
@@
According to Eqs. (9) and (10), one can find that the variation of both population $p_{k}$ and coherence magnitude $C_{kk^{\prime}}$ are determined by the coefficient $\lambda^{(i)}_{k}(\tau)$. It is noted that $\lambda^{(i)}_{k}(\tau)$ is independent on the resonator state, and yet depends on the evolution period $\tau$, the resonator-qubit detuning $\delta$, and the resonator-qubit coupling strength $g$. By optimizing the evolution period $\tau$ for each round of evolution and measurement, one can construct a filter preserving the desired elements in the density matrix of the resonator with $|\lambda^{(i)}_{k}(\tau)|=1$ and in the same time eliminating the others with $|\lambda^{(i)}_{k}(\tau)|<1$.

### v2

# Generating Fock-state superpositions from coherent states by selective measurement

@@ Abstract
Fock states and their superpositions are exotic testbeds for nonclassical physics and valuable resources for quantum technologies. We provide a simple protocol for the quantum measurement to generate an arbitrary Fock state and certain superposed Fock states from a coherent state of a target resonator, without any carefully tailored driving. This conditional protocol can be efficiently constructed by a sequence of joint free evolution of the resonator and an ancillary qubit, which are coupled via a Jaynes-Cummings interaction, and projective measurements on the qubit. By properly choosing the duration of each evolution-measurement cycle and the initial state of the resonator, we can generate a desired Fock state $|n\rangle$ and a superposed Fock state $(|0\rangle+|n\rangle)/\sqrt{2}$, $n\sim 10$, with a fidelity over $99\%$ in less than $30$ measurements. Moreover, our protocol can be extended straightforwardly to the generation of a Bell-like state $(|00\rangle+|nn\rangle)/\sqrt{2}$ with multiple excitations in a double-resonator system. We also calculate the outcome fidelity and the success probability of our protocol in the presence of decoherence.
@@ I Introduction
A Fock state is one of the most nonclassical states with a well-defined number of bosonic excitations such as photons, phonons, and magnons. It plays diverse and crucial roles in quantum information fields Yamamoto and Haus (1986), including quantum communications Caves and Drummond (1994), quantum metrology Holland and Burnett (1993); Nagata et al. (2007); Giovannetti et al. (2011); Slussarenko et al. (2017), and remote entanglement generation Narla et al. (2016). Fock states are prominent candidates in open quantum systems to explore the environment-induced decoherence by measuring their decay time Wang et al. (2008); Brune et al. (2008). The superpositions of Fock states with particular components exhibit appealing features in fault-tolerant quantum information processing Joshi et al. (2021). For example, the superposed states with binomial coefficients for the Fock basis can be used in quantum error correction Michael et al. (2016). In addition, a high-precision measurement over weak classical signals can be achieved by using superposed Fock states Caves et al. (1980) due to the squeezing effect induced by quantum interference between Fock components Wódkiewicz et al. (1987). Generating desired Fock states and their superposition is therefore of significant interest to both the fundamental quantum physics and the applications in quantum technology.
@@
Early proposals to generate the Fock state employed the resonator and the ancillary atom in the Jaynes-Cummings (JC) model Larson and Mavrogordatos (2021), which are under full control in both initial states and coupling period. The atom serves as an intermediator between the resonator and the external driving field to pump quanta into the resonator by sequential half Rabi oscillations. The oscillation period depends on the current state of the resonator, by which the generation proposals demand elaborate external driving in the cavity QED setup Vogel et al. (1993); Law and Eberly (1996). Based on such a stepwise method, Fock states and their superpositions were then experimentally generated in coplanar waveguide resonators for a microwave photon number $n\leq 6$ with a fidelity over $90\%$ Hofheinz et al. (2008, 2009) and in a bulk acoustic-wave resonator for up to $n=7$ phonons with a fidelity over $26\%$ Chu et al. (2018). A Fock state with $n\leq 3$ can be generated in the transmon-cavity platform by dispersively coupling the transmon qubit to the photonic resonator and using a two-photon stimulated Raman adiabatic passage Premaratne et al. (2017). Two-level atomic systems are not necessary in the step-by-step idea. A proposal in Ref. Yanagimoto et al. (2019) was presented to generate the Fock state by the simultaneous adiabatic modulation over the driving frequency and intensity on a cavity with Kerr nonlinearity.
@@
Beyond the methods based on multiple Rabi oscillations or equivalent transitions, Uria $et~{}al$. considered in a cavity QED setup, a two-level atom resonantly coupled to the target cavity field initially prepared in a coherent state Uria et al. (2020). By numerical calculation, they found that after a preassigned period of free evolution, the cavity field evolves into a displaced Fock state of a large photon number with a fidelity over $70\%$. Recently, the experimental demonstration of a dissipation engineering over a three-dimensional microwave cavity was reported in Ref. Li et al. (2024), which is based on a cascaded selective photon-addition operation assisted by an ancillary superconducting qubit and a readout resonator. The readout resonator was used to implement a quantum reservoir control for stabilizing a multiphoton Fock state of the cavity mode with a photon number up to $n=3$ with a fidelity over $77\%$.
@@
Quantum measurement is found to be a promising candidate for state control tasks, such as quantum state purification Nakazato et al. (2003); Combes et al. (2010); Wiseman and Ralph (2006), ground-state cooling of various quantum systems Li et al. (2011); Puebla et al. (2020); Yan and Jing (2021), and nuclear spin polarization Jin et al. (2022). Also, it can be implemented in generating Fock states by elaborate parametric tuning or by nondemolition measurements. For example, a Fock state of a cavity field can be conditionally created by the postselection over the atomic level after a specified period of evolution in a cavity QED setup, in which the atom-field interaction is highly tunable to control the transition rate in a chosen atom-field subspace on resonance Fran ça Santos et al. (2001). An arbitrary superposed Fock state can be generated by the dispersive coupling between the cavity field and a qubit so that one could individually control the qubit transition rate associated with the corresponding qubit-field subspace by driving the qubit with external microwave signals containing multi frequency components Wang et al. (2017). On performing quantum nondemolition measurement and feedback control over the field photon number to conditionally suppress the photon-number spread in Fock space Guerlin et al. (2007), a Fock state with a small photon number $n\leq 3$ was proposed in theory Dotsenko et al. (2009) and implemented in experiment Sayrin et al. (2011). Recently, quantum nondemolition measurement has been used in the preparation of a superposed Fock state in a circuit QED setup, which can be optimized by reinforcement learning Perret and Bérubé-Lauzière (2024).
@@
With a JC Hamiltonian, we propose a protocol to conditionally generate a desired Fock state $|n\rangle$ and certain superposed Fock states, such as $(|0\rangle\pm|n\rangle)/\sqrt{2}$, of a target resonator that starts from the coherent state. The protocol is based on repeated projective measurements on the coupled qubit, which are separated with joint free evolution of the composite system lasting $\tau$. In the near-resonant case, $\tau$ is much shorter than that under the dispersive coupling. For the resonator, the evolution-and-measurement cycles induce a positive operator valued measure (POVM) described by $V_{ji}(\tau)=\langle j|e^{-iH\tau}|i\rangle$, where $|i\rangle$ and $|j\rangle$ represent the initial state of the ancillary qubit and the measurement outcome, respectively and $H$ denotes the full Hamiltonian of the composite system. Our protocol does not require any elaborate external driving or parametric tuning during the whole generation process. By appropriately choosing a POVM and the period $\tau$ for each evolution-measurement cycle, the selected Fock components can be gradually distinguished from the coherent state. By using a qutrit instead of a qubit and two nondegenerate resonators instead of a single one, our protocol can be extended to generate Bell-like states $(|00\rangle\pm|nn\rangle)/\sqrt{2}$. In the presence of decoherence, the outcome fidelity of our protocol is found to be limited by the trade-off between the projective measurements and the decay from the desired Fock components to their lower neighbors. In terms of the JC model and the projective measurement over the ancillary system, our work is inspired by previous proposals for generating nonclassical bosonic states, e.g., the Fock states Harel and Kurizki (1996), the classical mixture of Fock states Delakouras et al. (2023), and the squeezed state in Ref. Weiher et al. (2019). Nevertheless, our protocol is applicable to the generation of superposed states and extendable to multiple modes.
@@
The rest of this work is organized as follows. In Sec. II we introduce the theoretical framework about the POVM based on the JC model and repeated projective measurements on the ancillary qubit. We find the reduction factor of the population distribution of the resonator, which is crucial to the ensuing state generation and is determined by the qubit-resonator coupling strength, their detuning, and the period of each evolution-and-measurement cycle. In Sec. III.1 and Sec. III.2 we discuss the generation of various single-mode Fock states and their coherent superpositions, respectively. The protocol efficiency can be optimized by only modulating the evolution periods. In Sec. IV our protocol is extended to generate Bell-like state of multiple excitations in two nondegenerate resonator modes, which can be described by the two-dimensional population-reduction factors. We summarize the whole work in Sec. V.
@@ II Theoretical framework
Figure 1: Model diagram for our superposed Fock-state-generation protocol. A two-level system (qubit) serves as the ancillary system. The target resonator is coupled to the qubit through exchange interaction with strength $g$ and $\delta$ their detuning in frequency. The projective measurements $M_{i}=|i\rangle\langle i|$, $i=g,e$, are performed on the ancillary qubit.
To generate the desired Fock states or their superposition of a single bosonic mode, our protocol is based on a composite system consisting of a target resonator coupled to an ancillary qubit with the JC interaction Larson and Mavrogordatos (2021), as shown in Fig. 1. The full Hamiltonian reads ($\hbar=1$)
@@
where $\omega_{e}$ and $\omega_{f}$ are the level splitting of the ancillary qubit and the transition frequency of the target resonator, respectively, $a^{\dagger}$ ($a$) is the creation (annihilation) operator of the resonator, and $g$ denotes the qubit-resonator coupling strength. In the rotating frame with respect to $U_{0}=\exp[i\omega_{f}(|e\rangle\langle e|+a^{\dagger}a)t]$, the Hamiltonian becomes
@@
where $\delta\equiv\omega_{e}-\omega_{f}$ denotes the detuning between the qubit and the resonator mode. Our theoretical framework is primarily established on the close-system Hamiltonian (2) to distinguish the role played by projective measurement. The decoherence effect will be discussed in Secs. III.1 and IV.
Suppose that the ancillary qubit is prepared as one of its energy eigenstates $|i\rangle$, $i\in\{e,g\}$, and the resonator is initially at $\rho_{a}(0)$. After a period $\tau$ of joint free evolution governed by $U(\tau)=\exp(-iH^{\prime}\tau)$, we perform a projective measurement on the qubit. If the measurement outcome is $|j\rangle$, $j\in\{e,g\}$, the density matrix of the resonator is then found to be
@@
where $V_{ji}(\tau)=\langle j|e^{-iH^{\prime}\tau}|i\rangle$ denotes an effective (nonunitary) evolution operator acting on the resonator and $P_{j}(1)$ represents the success probability of measuring the ancillary qubit on the state $|j\rangle$ after one evolution-measurement cycle. In our protocol, the cycle is repeated if $|j\rangle=|i\rangle$. Otherwise, the entire protocol is restarted from the very beginning.
The effective evolution operator $V_{ii}(\tau)$ then can be relabeled as $V_{i}(\tau)$ for simplicity, which is diagonal in the Fock-state basis $\{|k\rangle\}$ Li et al. (2011); Yan and Jing (2021), i.e.,
$V_{i}(\tau)=\sum_{k=0}^{\infty}\lambda_{k}^{(i)}(\tau)|k\rangle\langle k|.$ (4)
@@
where the $k$-photon Rabi frequencies for $i=e,g$ are
@@
respectively. Since $|\lambda^{(i)}_{k}(\tau)|\leq 1$, it works evidently as a reduction factor for the population on the Fock-state $|k\rangle$.
If the outcomes of each measurement are identical to the initial state of the qubit $|i\rangle$ and $\tau$ is invariant, the density matrix of the resonator for $N$ rounds of evolution and measurement is found to be
$\rho_{a}(N\tau)$ $=\dfrac{V_{i}^{N}(\tau)\rho_{a}(0)V_{i}^{\dagger N}(\tau)}{\text{Tr}\left[V_{i}^{N}(\tau)\rho_{a}(0)V_{i}^{\dagger N}(\tau)\right]}$
@@
$\mathcal{D}[\rho_{a}(N\tau)]$ $=\sum_{k=0}^{\infty}\frac{p_{k}|\lambda_{k}^{(i)}(\tau)|^{2N}}{P_{i}(N)}|k\rangle\langle k|,$ (9)
$\mathcal{C}[\rho_{a}(N\tau)]$ $=\sum_{k\neq k^{\prime}}\frac{C_{kk^{\prime}}\left[\lambda_{k}^{(i)}(\tau)\lambda_{k^{\prime}}^{(i)*}(\tau)\right]^{N}}{P_{i}(N)}|k\rangle\langle k^{\prime}|,$ (10)
where $\mathcal{D}[\rho_{a}]$ and $\mathcal{C}[\rho_{a}]$ denote the diagonal and off-diagonal parts of the density matrix, respectively. And $p_{k}\equiv\langle k|\rho_{a}(0)|k\rangle$ and $C_{kk^{\prime}}\equiv\langle k|\rho_{a}(0)|k^{\prime}\rangle$ are the populations and coherence of the initial density matrix of the resonator in the Fock-state basis $\{|k\rangle\}$. The success probability of finding the qubit at its initial state after $N$ measurements is found to be
@@
According to Eqs. (9) and (10), one can find that the variations of both the population $p_{k}$ and coherence magnitude $C_{kk^{\prime}}$ are determined by the coefficient $\lambda^{(i)}_{k}(\tau)$. Note that $\lambda^{(i)}_{k}(\tau)$ is independent of the resonator state, and yet depends on the evolution period $\tau$, the resonator-qubit detuning $\delta$, and the resonator-qubit coupling strength $g$. By optimizing the evolution period $\tau$ for each round of evolution and measurement, one can construct a filter preserving the desired elements in the density matrix of the resonator with $|\lambda^{(i)}_{k}(\tau)|=1$ and in the same time eliminating the others with $|\lambda^{(i)}_{k}(\tau)|<1$. We stress that the filter function in our protocol can be used to construct the superposed Fock state without parametric modulation over system frequencies and coupling strength Larson and Stenholm (2004).